OpenStudy (anonymous):

Which of the following is the solution of log (x+3) 0.001 = -3 ? x = 5 x = 7 x = 13 x = 15

OpenStudy (anonymous):

x+3 is the base.

OpenStudy (anonymous):

\[\log_{x+3} 0.001=-3\] since 0.001=10^-3 so,\[\log_{x+3} 10^{-3} = -3\] \[-3\log_{x+3} 10 = -3\] \[\log_{x+3} 10=1\] \[(\log10)/(log(x+3))=1\] \[\log 10=\log(x+3)\] hence, 10=x+3, x=7

OpenStudy (precal):